Arcadia HS, purchased equipment for $510,000 which was estimated to have a useful life of 10 years with a residual value of $10,000 at the end of that time. Depreciation has been recorded for 7 years on a straight-line basis. In 2014 (year 8), it is determined that the total estimated life should be 15 years with a residual value of $5,000 at the end of that time.

What is the journal entry to correct

the prior years’ depreciation?

Calculate the depreciation expense

for 2014.Arcadia HS, purchased equipment for $510,000 which was estimated to have a useful life of 10 years with a residual value of $
Answer #1

b) Accumulated depreciation = (510000-10000)/10*7 = 350000

Revised depreciation = (510000-350000-5000)/8 = 19375